The base-dissociation constant of ethylamine (C_2H
Answer #1

      C2H5NH2 + H2O ---------> C2H5NH3+ + OH-

I              0.016                                  0               0

C             -x                                       +x              +x

E           0.016-x                                  +x             +x

                    Kb = [C2H5NH3+][OH-]/[C2H5NH2]

                     6.4*10-4   = x*x/0.016-x

                     6.4*10-4 *(0.16-x) = x2

                            x= 0.0098

                 [OH-]   =x= 0.0098M

                Kw      = [H+][OH-]

                 [H+]   = Kw/[OH-]

                        = 1*10-14/0.0098

                        = 1.02*10-12 M >>>> answer